OpenStudy (anonymous):

Tena ate one of the muffins she baked and had 10 left. How many muffins did Tena bake? Write an equation. Then solve the problem.

OpenStudy (anonymous):

a. x + 1= 10; x=11 c. x - 1= 10; x=11 b. x +1= 10; x=10 d. x - 1= 10; x=10

OpenStudy (anonymous):

c. because one muffin is taken away, and the total amount of muffins (x) is 11

OpenStudy (anonymous):

It is definitely C

OpenStudy (anonymous):

thanks!!